OpenStudy (anonymous):

Prove: Cos θ - Sin θ/Cosθ + Sin θ = Cot θ-1/Cot θ+1

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

Yes please @OOOPS

OpenStudy (anonymous):

divided both numerator and denominator of the left hand side by sin \(\theta\), just one line you get the answer, hehehe...

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Ohh... Thank you @OOOPS

OpenStudy (anonymous):

easy, right?

OpenStudy (anonymous):

hey, if you don't get, just speak out. I can illustrate what I say

OpenStudy (anonymous):

Sadly I was absent one day and I get a worksheet with 19 questions and I need to choose 15 of them... No one bothered to explain anything so yeah I'm lost... @OOOPS

OpenStudy (anonymous):

ok, let \(\theta\) aside, just indicate the trig, if I say sin, that means sin\(\theta\), if I say cot, that means cot \(\theta\), ok?

OpenStudy (anonymous):

reply to show you understand me, please

OpenStudy (anonymous):

I understand that yes

OpenStudy (anonymous):